OpenStudy (anonymous):

Any nerds want to help out? How to you graph equations like for example: y=2x-7 Equation 1 y=-4x+5 Equation 2 Thank you so much!!!

jimthompson5910 (jim_thompson5910):

To graph any straight line, you just need two points To get any point, you plug in any x value to get a corresponding y value For instance, if x = 0, then y = 2x - 7 y = 2*0 - 7 y = 0 - 7 y = -7 which tells us that the point (0, -7) is on the line y = 2x - 7

OpenStudy (anonymous):

Nah I'm not a nerd. I'm a gangster. I'm also a bike junkie. I also smoke a pack a day.

jimthompson5910 (jim_thompson5910):

you repeat that process for some other x value (say x = 1) to get another point Once you have your two points, plot them and draw a straight line through them.

OpenStudy (anonymous):

Confusing but Thanks:0)

jimthompson5910 (jim_thompson5910):

where are you lost

OpenStudy (anonymous):

It just doesn't seem like what we did in math

jimthompson5910 (jim_thompson5910):

how did you graph? there are multiple ways to do it you could start with the y-intercept, then use the slope to find another point

OpenStudy (anonymous):

plot points.

jimthompson5910 (jim_thompson5910):

to graph y = 2x - 7, you start with the y-intercept (0, -7) then you go up 2 and over to the right 1 to get the next point (1,-5) so you now have two points to draw a straight line through

OpenStudy (anonymous):

The section in my math book is titled Solve Linear Systems by Graphing

OpenStudy (anonymous):

So is what is 7? the y?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

do you see how I'm getting this though? hopefully you see how to graph any linear equation

jimthompson5910 (jim_thompson5910):

the -7 is the y-intercept, which is the point (0, -7)

OpenStudy (anonymous):

@jim_thompson5910 nailed it on the head.

OpenStudy (anonymous):

Okay so why is there still a y in the equation if it is -7

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

Forget the question...:)

jimthompson5910 (jim_thompson5910):

you sure?

jimthompson5910 (jim_thompson5910):

what I did when i plugged in x = 0 to get y = -7 shows that we have the point (0, -7) on the line y = 2x - 7

OpenStudy (anonymous):

thanks for your help!!!! greatly appreciated....

jimthompson5910 (jim_thompson5910):

ok yw